You have a mean score of 84 after taking five 100-point tests. What do you need to score on the sixth 100-point test to have a mean score of 85?
I tried: 84+x/2=85
2(84+x/2)=(85)(2)
84-84+x=170-84
x=86
The problem is that I have multiple choice answers and 86 is not one of the answers (the answers are: A. 85 B. 89 C. 90 D. 92

Answer by ewatrrr(24785)


Hi,
Let x represent the score on the sixth test
Mean = sum of scores/number of tests
84 = S/5 test Therfore 5*84 = 420 is the sum of the first five tests
(420 + x)/6 = 85
420 + x = 510
x = 90